# A neural networks method to search for long transient gravitational waves

### v2

@@ I Introduction
Several other source classes are searched in GW data [6], and have not been detected yet. Among them, there are sources of continuous quasi-periodic GWs (CW), e.g., asymmetric spinning neutron stars [7]. For such sources, the signal frequency is linked to the star’s rotational frequency by a proportionality factor and slowly decreases in time at a rate (i.e. spin-down) $\lesssim(10^{-10}\mathrm{Hz}/\mathrm{s})$, due to the emission of GW itself and likely other energy losses, like the emission of electromagnetic (EM) radiation. So far, interesting upper limits on the amplitude of a possible CW signal have been placed (see e.g., [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]).
@@
A variety of models (e.g. [20, 21, 22, 23]) have been proposed for millisecond spinning, highly asymmetric NS (likely, newborn magnetars) emitting quasi-periodic GW signals over a limited time interval (in the detector’s band), from minutes to several hours, with a spin-down $\sim\mathcal{O}(\mathrm{Hz}/\mathrm{s})$. Such signals are generically called long transients (tCW) [24]. As we currently do not know exactly how matter behaves at supra-nuclear densities, the detection of this kind of signals will be a powerful additional tool in the study of the NS interior structure and behavior. Typical search techniques used for tCWs, based on semi-coherent methods (see e.g., [25, 26, 27, 28], applied for istance in [29, 30]) or on matched filter (see e.g., [31, 32]), are computationally demanding. Here we propose a machine learning (ML) technique to perform a robust and computationally feasible search and produce a first list of candidates. With the exception of the initial studies done in [33, 34] there are no ML procedures implemented for this kind of signal. In general, data analysis using neural networks (NNs) requires less computing power than more classical algorithms. Indeed, once a model [^1] Model aka neural network in this paper. is trained, it can be applied to different cases, as we will see in the subsection IV.5. The aim of this work is to develop ML techniques to analyze ground-based interferometric detector data. This paper presents the implementation of a NN classifier capable of identifying the presence or absence of tCW signals. Data are represented using time-frequency maps, input of the classifier model built with convolutional neural networks (CNN) [35]. An additional neural network model has been placed before the classifier itself to improve its performance, specifically a NN denoiser model exploiting residual learning [36].
@@
The paper is organized as follows. In Sec. II, we describe the main features of the tCW expected from a newly born magnetar. In Sec. III we explain how we build time-frequency maps, input to our artificial neural networks, and how we choose the parameter space to simulate the signal. In Sec. IV, we report the results of our study, describing first the denoiser and classifier performances and subsequently the robustness of our models. Finally, we draw conclusions in Sec. (V).
@@ II Signal model
In general, an isolated NS spinning with frequency $f_{\rm rot}$ can emit GWs if it presents an asymmetry with respect to the spin axis. GW emission due to the NS time-varying mass quadrupole corresponds to  [^2] Some instabilities, e.g. $r-$modes, lead to GW emission with $n=7$. $n=5$ (e.g., [38]). If, in addition, the rotational energy is radiated only via GWs, then the spin-down equation (1) becomes
@@
$\epsilon\equiv\frac{I_{1}-I_{2}}{I},$ (3)
@@
We trained our model assuming $n=5$, that is, without taking into account the electromagnetic contribution to the spin-down (which is described, for instance, in [45]). To assess the robustness of the model, we tested its performance by injecting signals with different braking indexes between 3.5 and 5, as discussed in Sec. (IV). Our goal in the future is to generalize our model by combining electromagnetic and gravitational torques.
@@ III Methods
III.1 Time-frequency maps
@@
Figure 1: Example of spectrograms of noise plus signal, only signal and only noise. The physical parameters for the signal are $f_{0}$=1982 Hz, $\epsilon$=0.0028, $h_{0}=2\times 10^{-23}$. Each map has a frequency range between 1850 Hz and 2000 Hz and covers a time interval of 1200 s. The noise is simulated according to the O4555Begin: May 24, 2023, end: planned February 2025. design sensitivity curve of LIGO interferometers.
Fig. (5) shows examples of spectrograms, with time on the horizontal axis and frequency on the vertical axis, computed from simulated colored noise [48] following the LIGO O4 design sensitivity curve [49]. The color bar corresponds to amplitude spectral density (ASD), i.e. the square root of the spectrum, associated with each pair of time and frequency values.
@@
The choice of the time bin $\delta t$ is based on the characteristics of the signals. Ideally, in order to maximize the signal-to-noise ratio, one would choose the longest duration, which still allows the signal power to remain confined in a single frequency bin, i.e. ($\dot{f}\delta t\leq\delta f=1/(2\delta t)$). However, different signals can have widely different values of $\dot{f}$, i.e. different slopes in the time-frequency plane, depending on the signal parameters. Given the parameter space we explore (see Eq. (12) and related discussion), after various tests, we have concluded that a reasonable choice is to use a segment duration $\delta t$ = 2 seconds. In addition, in order to build squared maps, we computed each spectrogram over a time interval of 1200 s and a frequency band of 150 Hz. In this way, we have maps of 600x600 pixels, where each pixel has a single value assigned. So a time-frequency map can be represented as a gray scale image. The choice of this representation was suggested by the properties of the signal we are searching for, as discussed later in this section. The size (600x600) was also chosen, taking into account the maximum memory that can be used on the GPU in NN training and the layers we used to implement the model.
Since we are computing the FFT of segments of finite duration, we have to account for spectral leakage, which consists in a spread of the power content of a frequency bin among neighboring bins. To reduce spectral leakage, we adopted a flat-top window, defined as a sum of cosine functions:
$f_{\rm cos}(x_{i})=\sum_{k=0}^{4}(-1)^{k}a_{k}\cos\bigg{(}\frac{2k\pi x_{i}}{N}\bigg{)},$ (11)
@@
To train the NN needed for the analysis, we created three types of maps containing, respectively, only noise, only the signal and the sum of noise plus signal, as shown in Fig. (5), from left to right. To simulate the signal, according to Eq. (6), we fixed a fiducial value for the NS momentum of inertia, $I=1.4\times 10^{38}\ \mathrm{Kg\ m}^{2}$. As for the angle of sight, we used $\iota=55^{\circ}$, i.e. the mean value for a uniform distribution of $\cos{\iota}$. Finally, we set the following ranges for the initial frequency ($f_{0}$) and NS $\epsilon$, respectively:
@@
As shown in Eq. (4), different values of $f_{0}$ and $\epsilon$ lead to widely different signal shapes in the time-frequency plane. In particular, a higher initial frequency and/or larger ellipticity imply a much faster frequency evolution, which, in some cases, may decrease by over 150 Hz in less than $1200$ s. Such signals would therefore cross multiple maps; to augment the chances of retrieving them maps were interlaced by 75 Hz. As an example, Fig. (2) depicts a signal with $f_{0}=1985\ \mathrm{Hz}$ and $\epsilon=2.8\times 10^{-3}$ covering a range wider than 300 Hz in 1200 s, thus appearing in 6 interlaced maps.
@@
III.2 Machine learning models
@@
III.2.1 Denoiser
The goal of a denoiser, a specialized ANN, is to retrieve a clean image $y$ from a noisy observation $x=y+\nu$, where $\nu$ is the noise. We adopted the residual learning approach (see [55] and appendix (A) for details on our modified architecture). The model is closed by a skip connection that links the output to the input. The idea is to teach the model how to predict noise and then subtract it from noisy data images in order to produce a denoised map, i.e. the input for the classifier.
@@ III.2.2 Classifier
The receptive field in the first layer, as shown in appendix (B), is bigger for the classifier than for the denoiser because we want to identify a correlation between distant pixels on the scale of the whole map. The signal differs from the noise because of this correlation, so this is a key feature that we want to extract to classify the maps.
@@
We studied the performance of the classifier, in particular the efficiency of detection (Eff) and the false alarm probability (FAP), adopting different probability thresholds to discriminate between the presence or absence of a signal in a given map. In particular, we computed the FAP as
@@
where the denominator is the sum of all the samples considered. Eventually, to obtain a good balance between Eff and FAP, we chose the F1 score, [59], defined as:
@@
F1 is a harmonic mean, so it gives more weight to lower values of P and R, and it is large only if both P and R are large. In particular, it is 0 if there are no true positives and 1 if there are no false predictions.
@@ IV.1 Dataset preparation
To train the denoiser, we simulated 1000 signals in the range described by Eq. (12) and 200 more with initial frequency $f_{0}$ in the frequency interval [1.8, 2.0] kHz. To test the denoiser and later train and test the classifier, we built another set of maps, with 2400 signals in the same intervals as Eq. (12). We trained and tested our models for different initial amplitudes, and we report here the results for $h_{0}(t_{0})=2\times 10^{-23}$.
@@
The two plots in Fig. (3) summarize the main properties of the denoiser training set. On the left panel, we report the number of maps crossed by each signal, and on the right panel, the $pr_{sn}$ (Eq.13) of each signal versus the NS ellipticity ($\epsilon$) and initial spin frequency ($f_{0}$). Each point represents a different signal. Notice that the higher the initial frequency and the ellipticity, the higher the number of maps crossed by the signal. Indeed, as described in Eq. (4), the frequency in this case varies more rapidly. As a result, the signal loses energy faster, and $pr_{\rm sn}$ is lower (see the upper right corner of the parameter space). As discussed in the next subsection, the region where both $f_{0}$ and $\epsilon$ are large is where both denoising and classification become more challenging. The test set exhibits the same trend.
@@ IV.2 Denoising
We trained the denoiser as described in Sec. (III.2.1), and then evaluated the model’s performance in reducing the noise and in preserving the signal separately. In particular, while the training was done with ”noise plus signal” maps, the tests were carried out separately on ”only noise” maps and on maps containing both noise and a signal.
We first focus on the noise reduction capability and summarize our results in Fig. (4). The histograms on the left panel show the mean value of all pixels in each map of the test set (”noise only” maps) before ($\mu_{\mathrm{noise}}$, orange) and after ($\mu_{\mathrm{denoised}}$, blue) the denoiser. On the right panel, we show the ratio of the means, defined as
$\mu_{\mathrm{ratio}}=\frac{\mu_{\mathrm{noise}}}{\mu_{\mathrm{denoised}}}\,,$ (19)
@@
We then turned to evaluating the effectiveness of the denoiser in preserving the signal by testing it on ”noise+signal” maps. Given that, as discussed above, individual signals may appear in multiple maps, we report in Fig. (5) the best value of the overlap, $\mathcal{O}$ (Eq. 15), in the $f_{0}-\epsilon$ plane, i.e., the value of the overlap computed on the map where the signal has been denoised better. Indeed, we stress that, in order to have a valid ”signal trigger,” it is sufficient to be able to tag right one map. We note that there is a correlation between the parameters of the signals and the ability of the denoiser to retrieve them. Indeed, the worst values of overlap are found in the upper right corner, the same region where $pr_{\rm sn}$ is lowest (right panel of Fig. (3)), and the signal reaches the noise level more rapidly.
@@ IV.3 Classification
Our goal is to distinguish maps where the signal is present from those where it is absent. The training procedure is applied to our classifier, described in Sec. (III.2.2), both on maps that were passed through the denoiser (clean maps) and maps that were not passed through the denoiser (noisy maps). In the latter case, the training of the classifier fails (for an initial amplitude of 2E-23), i.e., the model does not find a set of parameters that minimize the training and validation losses. As a consequence, the model does not learn the difference between maps with signals and maps that contain only noise. This demonstrates the crucial role of the denoiser in enhancing the chances of detection. We thus went on to only classify clean maps, using the F1 score (Eq. 17) to find the best compromise between the false alarm probability (FAP) and the efficiency (Eff).
As discussed in the previous section, a signal can cross multiple maps when its frequency varies very rapidly. We define a ”positive trigger” to select a GW candidate when at least one map is correctly tagged. So, from here on, we will report the efficiency per signal (Eff), defined as the number of positive triggers over the total number of injected signals.
@@
In Fig. (6), we show the FAP, Eff and F1 score for different values of the probability threshold of the classifier. The fuchsia straight line is the value that maximizes the F1 score and corresponds to
@@
where $\mathrm{FAP}_{1800}$ is the FAP computed for noise-only maps whose frequency range is under 1800 Hz (blue crosses) and $\mathrm{Eff}_{1800}$ is the efficiency computed on signals with $f_{0}<1800$ (red crosses). It is interesting to note that $\mathrm{Eff}_{1800}$ is higher than Eff and decreases slower.  Indeed, as shown in Fig. (5), as the frequency increases, it is more difficult to retrieve the signal. On the other hand, the FAP is not sensitive to the frequency and does not differ from $\mathrm{FAP}_{1800}$.

### v4

# A neural network method to search for long transient gravitational waves

@@ I Introduction
Several other source classes are searched for in GW data [6], and have not been detected yet. Among them, there are sources of continuous quasi-periodic GWs (CW), e.g., asymmetric spinning neutron stars [7]. For such sources, the signal frequency is linked to the star’s rotational frequency by a proportionality factor and slowly decreases in time at a rate (i.e. spin-down) $\lesssim(10^{-10}\mathrm{Hz}/\mathrm{s})$, due to the emission of GW itself and likely other energy losses, like the emission of electromagnetic (EM) radiation. So far, interesting upper limits on the amplitude of possible CW signals have been placed (see e.g., [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]).
@@
A variety of models (e.g. [20, 21, 22, 23]) have been proposed for millisecond spinning, highly asymmetric NSs (likely, newborn magnetars) emitting quasi-periodic GW signals over a limited time interval (in the detector’s band), from minutes to several hours, with a spin-down $\sim\mathcal{O}(\mathrm{Hz}/\mathrm{s})$. Such signals are generically called long transients (or transient continuous waves, tCWs)[24]. As we currently do not know exactly how matter behaves at supra-nuclear densities, the detection of this kind of signals will be a powerful additional tool in the study of the NS interior structure and behavior. Typical search techniques used for tCWs, based on semi-coherent methods (see e.g., [25, 26, 27, 28], applied for instance in [29, 30]) or on matched filter (see e.g., [31, 32]), are computationally demanding.
Here we propose a machine learning (ML) technique to perform a robust and computationally feasible search and produce a first list of candidates. With the exception of the initial studies done in [33, 34] there are no ML procedures implemented for this kind of signal. In general, data analysis using neural networks (NNs) requires less computing power than more classical algorithms. Indeed, once a model [^1] Model aka neural network in this paper. is trained, it can be applied to different cases, as we will see in the subsection IV.5.
The aim of this work is to develop ML techniques to analyze ground-based interferometric detector data. This paper presents the implementation of a NN classifier capable of identifying the presence or absence of tCW signals. Data are represented using time-frequency maps, inputs to the classifier model built with convolutional neural networks (CNN) [35]. An additional neural network model has been placed before the classifier itself to improve its performance, specifically a NN denoiser model exploiting residual learning [36].
@@
The paper is organized as follows. In Sec. II, we describe the main features of the tCW expected from a newly born magnetar. In Sec. III we explain how we build time-frequency maps, input to our artificial neural networks, and how we choose the parameter space to simulate the signal. In Sec. IV, we report the results of our study, describing first the denoiser and classifier performances and subsequently the robustness of our models. Finally, we draw conclusions in Sec. V.
@@ II Signal model
In general, an isolated NS spinning with frequency $f_{\rm rot}$ can emit GWs if it presents an asymmetry with respect to the spin axis. GW emission due to the NS time-varying mass quadrupole corresponds to $n=5$ (e.g., [38]), although some instabilities, e.g. $r-$modes, lead to GW emission with $n=7$. If, in addition, the rotational energy is radiated only via GWs, then the spin-down equation (1) becomes
@@
$\epsilon\equiv\frac{|I_{1}-I_{2}|}{I},$ (3)
@@ III Methods
The method we present involves the implementation of an NN that properly initialized allows the presence of a tGW to be distiguished from the noise present in the data produced by an interferometer. The data format we use and the NNs implemented are described here.
III.1 Introductory concepts of machine leaning
Machine learning is a type of Artificial Intelligence (AI) that allows computers to learn without being explicitly programmed [46]. It involves feeding data into algorithms (implemented in a network model) that can then identify patterns and make predictions on new data. ML implementations are classified into four major categories, depending on the nature of the learning “signal” or “response” available to a learning system. Among these we use Supervised Learning, that is the ML task of learning a function that maps an input to an output based on example input-output pairs. Both ”classification” and ”denoising” problems are supervised learning problems, solved with specific models (classifiers, denoisers) A classifier is a model that tries to predict the correct category between distinct classes(”labels”) of a given input data. The denoiser task involves removing noise from an image (from a data representation in which signal and noise is present). In our task, the models are fully initialized to use (”trained”) using the labeled training data(signal vs absence of signal). Before we get to the implementation of our models we introduce the data format we use in the next paragraph.
III.2 Time-frequency maps
@@
Figure 1: Example of spectrograms of noise plus signal, only signal and only noise. The physical parameters for the signal are $f_{0}$=1982 Hz, $\epsilon$=0.0028, $h_{0}=2\times 10^{-23}$. Each map has a frequency range between 1850 Hz and 2000 Hz and covers a time interval of 1200 s. The noise is simulated according to the O4 (begin: May 24, 2023, end: planned June 2025) design sensitivity curve of LIGO interferometers.
Figure 1 shows examples of spectrograms, with time on the horizontal axis and frequency on the vertical axis, computed from simulated colored noise [48] following the LIGO O4 design sensitivity curve [49]. The color bar corresponds to amplitude spectral density (ASD), i.e. the square root of the spectrum, associated with each pair of time and frequency values.
@@
The choice of the time bin $\delta t$ is based on the characteristics of the signals. Ideally, in order to maximize the signal-to-noise ratio, one would choose the longest duration, which still allows the signal power to remain confined in a single frequency bin, i.e. ($\dot{f}\delta t\leq\delta f=1/(2\delta t)$). However, different signals can have widely different values of $\dot{f}$, i.e. different slopes in the time-frequency plane, depending on the signal parameters. Given the parameter space we explore (see Eq. (12) and related discussion), after various tests, we have concluded that a reasonable choice is to use a segment duration $\delta t$ = 2 seconds. In addition, in order to build squared maps, we computed each spectrogram over a time interval of 1200 s and a frequency band of 150 Hz. In this way, we have maps of 600$\times$600 pixels, where each pixel has a single value assigned. So a time-frequency map can be represented as a gray-scale image. The choice of this representation was suggested by the properties of the signal we are searching for, as discussed later in this section. The size (600$\times$600) was also chosen also taking into account the maximum memory that can be used on the GPU in NN training and the layers we used to implement the model.
Since we are computing the FFT of segments of finite duration, we have to account for spectral leakage, which consists of a spread of the power content of a frequency bin among neighboring bins. To reduce spectral leakage, we adopted a flat-top window, defined as a sum of cosine functions:
$f_{\rm cos}(x_{i})=\frac{1}{2}\bigg{(}1-\cos\bigg{(}\frac{4\pi x_{i}}{N}\bigg{)}\bigg{)}\ \ \ \ i\leq\frac{N}{4},\ i\geq\frac{N}{4}$ (11)
$f_{\rm cos}(x_{i})=1\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{N}{4}<i<\frac{N}{4}$
@@
To train the NN needed for the analysis, we created three types of maps containing, respectively, only noise, only the signal and the sum of noise plus signal, as shown in Fig. 1, from left to right. To simulate the signal, according to Eq. (6), we fixed a fiducial value for the NS momentum of inertia, $I=1.4\times 10^{38}\ \mathrm{kg\ m}^{2}$. As for the angle of sight, we used $\iota=55^{\circ}$, representative of the average inclination. Finally, we set the following ranges for the initial frequency ($f_{0}$) and NS $\epsilon$, respectively:
@@
As shown in Eq. (4), different values of $f_{0}$ and $\epsilon$ lead to widely different signal shapes in the time-frequency plane. In particular, a higher initial frequency and/or larger ellipticity imply a much faster frequency evolution, which, in some cases, may decrease by over 150 Hz in less than $1200$ s. Such signals would therefore cross multiple maps. To augment the chances of retrieving them, maps were interlaced by 75 Hz. As an example, Fig. 2 depicts a signal with $f_{0}=1985\ \mathrm{Hz}$ and $\epsilon=2.8\times 10^{-3}$ covering a range wider than 300 Hz in 1200 s, thus appearing in 6 interlaced maps.
@@
III.3 Machine learning models
@@
III.3.1 Denoiser
The goal of a denoiser, a specialized ANN, is to retrieve a clean image $y$ from a noisy observation $x=y+\nu$, where $\nu$ is the noise. We adopted the residual learning approach (see [54] and Appendix A for details on our modified architecture). The model is closed by a skip connection [^3] Technique used to link network input data directly to the last stages by skipping some branches ([55]). that links the output to the input. The idea is to teach the model how to predict noise and then subtract it from noisy data images in order to produce a denoised map, i.e. the input for the classifier.
@@ III.3.2 Classifier
The receptive field in the first layer, as shown in Appendix B, is bigger for the classifier than for the denoiser because we want to identify a correlation between distant pixels on the scale of the whole map. The signal differs from the noise because of this correlation, so this is a key feature that we want to extract to classify the maps.
@@
We studied the performance of the classifier, in particular the efficiency of detection (Eff) and the false alarm probability (FAP), adopting different probability thresholds to discriminate between the presence or absence of a signal in a given map. In principle, the Eff should be computed as
$Eff=\frac{TP}{TP+FN}.$ (16)
Nevertheless, as discussed in subsection III.2, a signal can cross multiple maps when its frequency varies very rapidly. We define a ”positive trigger” to select a GW candidate when at least one map is correctly tagged. So, from here on, we will report the efficiency per signal (Eff), defined as the number of positive triggers over the total number of injected signals. For what it concerns the FAP, we computed it as
@@
where the denominator is the sum of all the samples considered. Eventually, to obtain a good balance between Eff and FAP, we used the F1 score, [59], defined as:
@@
F1 is a harmonic mean, so it gives more weight to lower values of P and R, and it is large only if both P and R are large. In particular, it is 0 if there are no true positives, i.e. Eff $\sim$ 0 and FAP $\sim$ 1, and 1 if there are no false predictions, i.e. Eff $\sim$ 1 and FAP $\sim$ 0. So we need to maximize the F1 score.
@@ IV.1 Dataset preparation
To train the denoiser, we simulated 1000 signals in the range described by Eq. (12) and 200 more with initial frequency $f_{0}$ in the frequency interval [1.8, 2.0] kHz. To test the denoiser and later train and test the classifier, we built another set of maps, with 2400 signals in the same intervals as Eq. (12). We trained and tested our models for different initial amplitudes, and we report here the results for $h_{0}(t_{0})=2\times 10^{-23}$. For more information, see the Appendix C.
@@
The two plots in Fig. 3 summarize the main properties of the denoiser training set. On the left panel, we report the number of maps crossed by each signal, and on the right panel, the $pr_{sn}$ (Eq.(13)) of each signal versus the NS ellipticity ($\epsilon$) and initial spin frequency ($f_{0}$). Each point represents a different signal. Notice that the higher the initial frequency and the ellipticity, the higher the number of maps crossed by the signal. Indeed, as described in Eq. (4), the frequency in this case varies more rapidly. As a result, the signal loses energy faster, and $pr_{\rm sn}$ is lower (see the upper right corner of the parameter space). As discussed in the next subsection, the region where both $f_{0}$ and $\epsilon$ are large is where both denoising and classification become more challenging. The test set exhibits the same trend.
@@ IV.2 Denoising
We trained the denoiser as described in Sec. III.3.1, and then evaluated the model’s performance in reducing the noise and in preserving the signal separately. In particular, while the training was done with ”noise plus signal” maps, the tests were carried out separately on ”only noise” maps and on maps containing both noise and a signal.
We first focus on the noise reduction capability and summarize our results in Fig 4 . The histograms on the left panel show the mean value of all pixels in each map of the test set (”noise only” maps) before ($\mu_{\mathrm{noise}}$, orange) and after ($\mu_{\mathrm{den}}$, blue) the denoiser, i.e.
$\mu_{\mathrm{noise}}=\sum_{t,f}\frac{1}{N_{pix}}N_{tf}^{\mathrm{noise}}\ \ \ \ \ \mu_{\mathrm{den}}=\sum_{t,f}\frac{1}{N_{pix}}N_{tf}^{\mathrm{den}}$ (20)
where $N_{tf}$ is the same as in Eq. (13). On the right panel, we show the ratio of the means, defined as
$\mu_{\mathrm{ratio}}=\frac{\mu_{\mathrm{noise}}}{\mu_{\mathrm{den}}}\,$ (21)
@@
We then turned to evaluating the effectiveness of the denoiser in preserving the signal by testing it on ”noise+signal” maps. Given that, as discussed above, individual signals may appear in multiple maps, we report in Fig. 5 the best value of the overlap, $\mathcal{O}$ (Eq. (15)), in the $f_{0}-\epsilon$ plane, i.e., the value of the overlap computed on the map where the signal has been denoised the best. Indeed, we stress that, in order to have a valid ”signal trigger,” it is sufficient to be able to tag right one map. We note that there is a correlation between the parameters of the signals and the ability of the denoiser to retrieve them. Indeed, the worst values of overlap are found in the upper right corner, the same region where $pr_{\rm sn}$ is lowest (right panel of Fig. 3), and the signal reaches the noise level more rapidly.
@@ IV.3 Classification
Our goal is to distinguish maps where the signal is present from those where it is absent. The training procedure is applied to our classifier, described in Sec. III.3.2, both on maps that were passed through the denoiser (clean maps) and maps that were not passed through the denoiser (noisy maps). In the latter case, the training of the classifier fails (for an initial amplitude of 2E-23), i.e., the model does not find a set of parameters that minimize the training and validation losses. As a consequence, the model does not learn the difference between maps with signals and maps that contain only noise. This demonstrates the crucial role of the denoiser in enhancing the chances of detection. We thus went on to only classify clean maps, using the F1 score (Eq. (18)) to find the best compromise between FAP and Eff.
@@
In Fig. 6, we show the FAP, Eff and F1 score for different values of the probability threshold of the classifier. The fuchsia straight line is the value that maximizes the F1 score and corresponds to
@@
where $\mathrm{FAP}_{1800}$ is the FAP computed for noise-only maps whose frequency range is under 1800 Hz (blue crosses) and $\mathrm{Eff}_{1800}$ is the efficiency computed on signals with $f_{0}<1800$ (red crosses). As we would expect, $\mathrm{Eff}_{1800}$ is higher than Eff and decreases slower. Indeed, as shown in Fig. 5, as the frequency increases, it is more difficult to retrieve the signal. On the other hand, the FAP is not sensitive to the frequency and does not differ from $\mathrm{FAP}_{1800}$.